OpenStudy (anonymous):

If a, b are natural numbers such that 2013 + a^2 = b^2, then the minimum possible value of ab is (A) 671 (B) 668 (C) 658 (D) 645

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Triangle_inequality Use the triangle inequality.

OpenStudy (anonymous):

But how to find the minimum value?

OpenStudy (ranga):

Let the product of a and b be called P. P = a * b ---- (1) 2013 + a^2 = b^2 ----- (2) Find b from (2), put in (1). You will have P as a function of just "a". Find dP/da, equate it to 0, solve for "a". Make sure it gives a minimum.

OpenStudy (ranga):

^^^ not working for me. :(

ganeshie8 (ganeshie8):

yeahh it gives a = b i think... which wont fit the equation

OpenStudy (anonymous):

The natural numbers part is throwing me off. Short of spamming 1 through sqrt(2013) until you find 2 natural numbers that fit, I'm drawing a blank.

OpenStudy (anonymous):

First think I notice is ab will grow when a or b grow. So the first natural number that makes 2013 + a^2 a square number will give a minimum product. Through trial and error, I found 14 and 47 work 14 * 47 = 658

OpenStudy (anonymous):

First *thing*

OpenStudy (anonymous):

This a list of some possible values of (a,b) {{14, 47}, {86, 97}, {334, 337}, {1006, 1007}}

OpenStudy (anonymous):

The min is 47(14)=658

OpenStudy (ranga):

Any other way to solve this problem besides trial and error?

OpenStudy (anonymous):

It is a Lagrange Multiplier problem Minimize f(a,b) = a b subject to \[ a^2 + 2013 = b^2 \]

OpenStudy (anonymous):

I tried Lagrange Multiplier. Didn't seem to work

OpenStudy (anonymous):

The answer is given as 658.Could u tell how to solve it by Lagrange Multiplier .

OpenStudy (anonymous):

I think Lagrange Multipliers does not seem to work with me.

OpenStudy (anonymous):

We do not have to forget that a and b are integers and not any real numbers. I wrote a mathematica script to generate my list

OpenStudy (anonymous):

I ran it from a=1 to 100000

OpenStudy (anonymous):

Here is the script Q = {}; For[a = 1, a <= 100000, a++, b = Sqrt[2013 + a^2]; If[ IntegerQ[ b] , Q = Append[Q, {a, b}]]] Q

OpenStudy (anonymous):

(b – -a) (b + a) = 2013 = 3 × 11 × 61 ab minimum when b -– a = 33 b + a = 61 a = 14

OpenStudy (anonymous):

But why b-a is minimum when it is 33,why not 3,11 or 61.

OpenStudy (ranga):

The possibilities are: (b-a) (b+a) 1 3x11x61 3 11 x 61 11 3 x 61 61 3 x 11 (not possible) 3x11 61 3x61 11 (not possible) 11x61 3 (not possible) 3x11x61 1 (not possible)

OpenStudy (anonymous):

you have have only 3 possibilities. 66 x 61 3 x 617 183 x 11 try each pair and see which one gives positive integer. As I said earlier, the first lowest integer will give the minimum product

OpenStudy (anonymous):

whoops i meant 33 x 61

OpenStudy (anonymous):

hold on, I messed up multiplicatoin

OpenStudy (anonymous):

it's 33 x 61 3 x 617 183 x 11

OpenStudy (anonymous):

(33 + 61)/2 = 47 (3 + 617)/2 = 310 (183 + 11)/2 = 97

OpenStudy (anonymous):

so the other solution is (47^2 - 2013)^.5 = 14

OpenStudy (anonymous):

Here is another thought, define \[ f(a )=a \sqrt{a^2+2013} \] This function is always increasing. At a=14 , f(14) = 658. a=14 is the first value that makes f(a) an integer. That does it

OpenStudy (anonymous):

it still involves trial and errors but at least it's still better than blindly use the calculator and change each natural number

OpenStudy (ranga):

The reason the earlier calculus methods did not work for this problem is because this is a discrete function. It is not continuous and cannot be differentiated to find the minima.

OpenStudy (anonymous):

This shows that f(14) is a minimum for integer values

OpenStudy (anonymous):

The discrete part is subset of the continuous version

OpenStudy (anonymous):

You only have to tests 14 numbers to get your result. That is not a lot even by hand.

ganeshie8 (ganeshie8):

Also, \(\large f(a )=a \sqrt{a^2+2013} \) forces \(\large a^2+2013\) to be a perfect square greater than 44 : 45/46/47

ganeshie8 (ganeshie8):

we get lucky in two trials for sqrt(a^2+2013) = 45, 46, 47... 47 works !

OpenStudy (anonymous):

Refer to the Mathematica attachement.